OpenStudy (anonymous):

Write an equation of a line that passes through the point (0, 4) and whose slope is 4. A. y = x + 4 B. y = 4x + 4 C. y = x – 4 D. y = x

OpenStudy (anonymous):

Equation of a line with slope and one point is given by: \[\large \color{green}{(y-y_1) = m(x-x_1)}\]

OpenStudy (australopithecus):

EQUATION OF THE LINE, MEMORIZE!!!! y = mx+b m = slope b = y-intercept we know slope so just plug into equation with the point

OpenStudy (anonymous):

Here m = 4 and \(x_1 = 0\), \(y_1 = 4\)

OpenStudy (australopithecus):

so we have y = 4x + b solve for b by subbing in the point into the equation, (0,4)

OpenStudy (australopithecus):

remember, (x, y)

OpenStudy (anonymous):

y=4x+ 4?

OpenStudy (australopithecus):

yes

OpenStudy (anonymous):

thank you